# Large language model psychometrics: A systematic review of evaluation, validation, and enhancement

### v2

@@ 3.2 Benchmarking with psychometrics-inspired principles
Task-oriented benchmarks often entail vast question sets to capture complex abilities. However, in many cases, the benchmarks either fail to fully represent these abilities due to their infinite manifestations or involve extraneous factors that are irrelevant to the target ability [Zhou et al., 2025, Wallach et al., 2025]. Recent research has drawn inspiration from psychometrics and explored the paradigm of construct-oriented evaluation, seeking its discriminative, predictive, and explanatory power. Ilić and Gignac [2024], Federiakin [2025] employ factor analysis to explore the latent variables underlying LLM benchmark performance. Their findings reveal a monolithic factor resembling general intelligence or ability. Federiakin [2025] ranks models based on this discovered factor and highlights its unique advantages over raw benchmark scores. In contrast, similar attempts by Burnell et al. [2023] identify three factors—reasoning, comprehension, and core language modeling—that better explain LLM performance across 27 cognitive tasks. Based on it, Zhu et al. [2024a] integrate the three factors into benchmark items to evaluate multifaceted abilities. This discrepancy between the estimated latent factors in the above findings can be attributed to differences in the models and benchmarks employed [Zhou et al., 2025]. Therefore, rather than relying on statistically derived factors, Zhou et al. [2025] propose a theory-driven hierarchical set of general scales for systematic construct-oriented evaluation. These scales are validated to explain what AI systems can do and predict their performance on novel task instances. Peng et al. [2024] present the Tong Test, a value- and ability-oriented framework, for Artificial General Intelligence (AGI) evaluation. This framework is rooted in dynamic embodied physical and social interactions (DEPSI), and can generate an infinite variety of tasks to evaluate key capabilities including values, learning, and cognition.
@@
Beyond defining and analyzing latent constructs, researchers have developed holistic, psychometrically rigorous methods for benchmark development. Liu et al. [2024e] introduce Evidence-Centered Benchmark Design (ECBD), a framework structuring benchmark creation into five modules—capability, content, adaptation, assembly, and evidence—each requiring justification to ensure validity. Through case studies of prominent LLM benchmarks, they demonstrate ECBD’s utility in identifying validity threats. Similarly grounded in psychometrics but with a distinct approach, Fang et al. [2024] propose Psychometrics-Assisted Benchmarking (PATCH), an eight-step process from construct definition to proficiency scoring. When piloted on 8th-grade mathematics, PATCH produced results diverging from traditional benchmarks, offering a more comprehensive evaluation. Building on related principles, Kardanova et al. [2024] adapt Evidence-Centered Design (ECD) to create psychometrically grounded benchmarks. Their application in pedagogy illustrates how this method can reduce data contamination and enhance test interpretability.
Item Response Theory.
Item Response Theory (IRT) is a statistical approach in psychometrics that jointly estimates the latent ability of examinees and the properties (difficulty, discrimination) of test items [Embretson and Reise, 2013]. Translating this framework to LLM evaluation enables researchers to infer latent ability scores, assess item informativeness, and perform more efficient evaluation. Recent research leverages principles from IRT to develop adaptive evaluation frameworks. These methods dynamically calibrate item difficulty based on model performance and weighting items by their inferred difficulty, aiming to achieve accurate evaluations with smaller test sizes and more discriminative items [Lalor et al., 2024, Polo et al., 2024, Zhuang et al., 2023b, a, Guinet et al., 2024]. Building on these approaches, Jiang et al. [2024a], Truong et al. [2025] introduce IRT-based benchmarks that involve learning to estimate item difficulty and learning to generate novel items calibrated to specific difficulty levels. Additionally, research has used IRT-based analyses to explore the alignment between LLM and human response distributions [He-Yueya et al., 2024]. IRT-based evaluations further offer the potential to estimate construct and item parameters on a unified scale, enabling direct comparisons across AI systems and against human norms, even when different test sets are used [Wang et al., 2023a, Fang et al., 2024].
@@ 4 Psychometrics for measuring psychological constructs
This section delves into the psychological constructs evaluated in LLM Psychometrics. Fig. 2 exemplifies the tests for the involved constructs.
Constructs Theories/inventories Related work
Personality traits Big Five Serapio-García et al. [2023], Jiang et al. [2023], Li et al. [2022b], Zou et al. [2024], Bhandari et al. [2025a], Huang et al. [2024c], Romero et al. [2024], Lee et al. [2024d], Liu et al. [2024a], Karra et al. [2022], Song et al. [2023], Caron and Srivastava [2022], Li et al. [2022a], Kovač et al. [2023], Frisch and Giulianelli [2024], Zheng et al. [2025], La Cava and Tagarelli [2024], Bodroža et al. [2024], Huang et al. [2023a], Gupta et al. [2024], Huang et al. [2023e], Pellert et al. [2024], Ai et al. [2024], Dorner et al. [2023], Klinkert et al. [2024], Petrov et al. [2024], Shu et al. [2024], Ren et al. [2024], Bhandari et al. [2025b], Hilliard et al. [2024], Jiang et al. [2024b], Lu et al. [2023], Sühr et al. [2023], Zhang [2024], Jain et al. [2024]
HEXACO Wang et al. [2025b], Bodroža et al. [2024], Miotto et al. [2022], Barua et al. [2024], Ren et al. [2024], Peereboom et al. [2024]
MBTI Rao et al. [2023], Pan and Zeng [2023], Song et al. [2024b], Cui et al. [2023], La Cava and Tagarelli [2024], Zhang et al. [2024a], Huang et al. [2023c], Ai et al. [2024], Lu et al. [2023], Chen et al. [2024a]
Dark Triad Li et al. [2022b], Romero et al. [2024], Lee et al. [2024d], Li et al. [2022a], Huang et al. [2023e], Barua et al. [2024], Peereboom et al. [2024], Lu et al. [2023]
Other & custom Wen et al. [2024b] (reviews); Mao et al. [2024a], Ai et al. [2024], Jiang et al. [2023], Zeng [2024]
Values Schwartz Pellert et al. [2024], Yao et al. [2025a], Shen et al. [2024], Kovač et al. [2024], Rozen et al. [2024], Zhang et al. [2023a], Yao et al. [2024], Hadar-Shoval et al. [2024], Fischer et al. [2023], Kovač et al. [2023], Miotto et al. [2022], Lee et al. [2024a], Ren et al. [2024], Ye et al. [2025a], Li et al. [2024b], Yao et al. [2025b], Cahyawijaya et al. [2024]
WVS Chiu et al. [2025], Yao et al. [2025a], Kim and Baek [2024], Faulborn et al. [2025]
VSM Kovač et al. [2023], Kharchenko et al. [2024], Zhong et al. [2024], Ren et al. [2024], Ye et al. [2025a]
GLOBE Li et al. [2024d], Karinshak et al. [2024], Ren et al. [2024]
SVO Zhang et al. [2024c]
Other & custom Zhang et al. [2024b], Li et al. [2024d], Biedma et al. [2024], Ye et al. [2025b], Meadows et al. [2024], Jiang et al. [2024a], Moore et al. [2024], Xu et al. [2023], Liu et al. [2024b]
Morality MFT Pellert et al. [2024], Nunes et al. [2024], Abdulhai et al. [2024], Neuman et al. [2025], Ji et al. [2024], Simmons [2023], Tlaie [2024], Aksoy [2024], Yan et al. [2024], Fraser et al. [2022], Zhou et al. [2024a], Yao et al. [2025b], Münker [2024]
ETHICS Jinnai [2024], Rodionov et al. [2023], Karpov et al. [2024], Yu et al. [2023], Albrecht et al. [2022]
DIT Khandelwal et al. [2024], Tanmay et al. [2023]
Other & custom Tanmay et al. [2023], Jin et al. [2022], Huang et al. [2024a], Jin et al. [2024b], Seror [2024], Bonagiri et al. [2024], Garcia et al. [2024], Ahmad and Takemoto [2024], Jiang et al. [2024a], Takemoto [2024], Han [2023], Peterson [2025], Ramezani and Xu [2023], Kucuk and Kocyigit [2023], Neuman et al. [2025], Chiu et al. [2025], Scherrer et al. [2023], Marraffini et al. [2024], Vida et al. [2024], Jinnai [2024], Meijer et al. [2024], Yuan et al. [2024], Liu et al. [2024d], Ohashi et al. [2024], Sachdeva and van Nuenen [2025]
Attitudes & opinions ANES Qi et al. [2024], Jiang et al. [2024d], Yang et al. [2024], Argyle et al. [2023], Bisbee et al. [2024], Sun et al. [2024], Jiang et al. [2022]
ATP Santurkar et al. [2023], Hwang et al. [2023], Tjuatja et al. [2024]
GLES Ma et al. [2024b], Ball et al. [2025], Qi et al. [2024], von der Heyde et al. [2024]
PCT Azzopardi and Moshfeghi [2024], Bernardelle et al. [2024], Wright et al. [2024], Röttger et al. [2024], Hartmann et al. [2023], Rozado [2023], Rutinowski et al. [2024]
Other & custom Ma et al. [2024a] (reviews); Dominguez-Olmedo et al. [2024], Kim and Lee [2023], Sanders et al. [2023], Kalinin [2023], Geng et al. [2024], Xu et al. [2025c], Lee et al. [2024c], Durmus et al. [2023], Rozado [2023], Rosenbusch et al. [2023], Wu et al. [2023], Chalkidis and Brandl [2024], Feng et al. [2023], Ceron et al. [2024], Röttger et al. [2025], Kim et al. [2025b], Faulborn et al. [2025]
Table 2: Representative LLM personality constructs in LLM psychometrics. The main dimensions or focus of the theories/inventories are listed in Table 3.
@@
MBTI Four dichotomies: Extraversion–Introversion, Sensing–Intuition, Thinking–Feeling, Judging–Perceiving
@@ 4.1 Measuring personality constructs
LLMs exhibit personality constructs that are not explicitly programmed or trained towards. The emergent constructs critically shape LLM behavior [Hagendorff, 2023], with profound implications for both individuals and broader social groups [Bengio et al., 2024]. Measuring these embedded psychological constructs is essential for understanding behavior, identifying biases, and fostering responsible development.
Table 2 and Table 3 summarize the representative personality constructs that have garnered attention in recent research. Researchers typically select constructs based on their relevance to LLM development and deployment, as well as the applicability of these constructs to AI systems. For instance, Li et al. [2024d] argue that emotional variability in LLMs is not a meaningful construct, given that LLMs lack the biological mechanisms underlying emotions. Conversely, personality and values are considered meaningful for LLMs, as they influence user interactions and model outputs [Serapio-García et al., 2023, Ye et al., 2025a]. This rationale has led to extensive research on personality, values, morality, and attitudes & opinions, as well as some—though less—focus on other constructs such as career selection [Hua et al., 2024], motivation [Chiu et al., 2025, Huang et al., 2023e], and mental health [Reuben et al., 2024, De Duro et al., 2024].
@@ 4.1.1 Personality traits
Personality is the enduring configuration of characteristics and behavior that comprises an individual’s unique adjustment to life [APA Dictionary of Psychology, ]. In the context of LLMs, personality traits relate to model safety, bias, and toxicity [Zhang et al., 2024a, Wang et al., 2025b], and highly determine user experience [Klinkert et al., 2024].
The study of personality traits leads to several prominent theoretical models, each offering unique insights into individual differences, including the Big Five [Goldberg, 2013], HEXACO [Ashton and Lee, 2007], Myers-Briggs Type Indicator (MBTI) [Myers et al., 1962], and Dark Triad [Paulhus and Williams, 2002]. When applying psychometrics to LLMs, most researchers directly administered the established inventories, such as NEO-PI-R [Costa and McCrae, 2008], BFI [John et al., 1991], and BFI-2 [Soto and John, 2017], for measuring the Big Five traits; HEXACO-60 [Ashton and Lee, 2009] and HEXACO-100 [Lee and Ashton, 2018] inventories for HEXACO traits; MBTI assessment [Myers, 1985] for MBTI types; and Dark Triad Dirty Dozen scale [Jonason and Webster, 2010] for Dark Triad traits.
Given widespread concerns about the practical relevance of these inventories [Ai et al., 2024], others have adapted them for more real-world scenarios. For example, Bhandari et al. [2025a] contextualize the tests in topic-specific or open-domain conversations; Ai et al. [2024] accompany the self-report tests with behavioral tests to examine the personality knowledge of LLMs; Jiang et al. [2023] compile and adapt existing tests into the Machine Personality Inventory; and Mao et al. [2024a] develop the PersonalityEdit inventory using LLMs. However, Peereboom et al. [2024] argue that human-derived traits may not meaningfully apply to LLMs, underscoring the necessity for psychometric theories specifically designed for LLM analysis.
@@
Early models such as GPT-3 score above human averages on Dark Triad scales [Li et al., 2022b, a, Romero et al., 2024]. Even after safety tuning, some models retain certain "dark qualities" [Li et al., 2022b], suggesting that these patterns may be deeply rooted in the pretraining data. More advanced LLMs indicate better level of alignment. They usually demonstrate high Openness, Extraversion, and Agreeableness traits, while low Neuroticism, on the Big Five personality tests [Zou et al., 2024, Bhandari et al., 2025a]. The results align with their design as assistive and helpful entities with emotional stability and engaging personality [Bhandari et al., 2025a, La Cava and Tagarelli, 2024]. Within the MBTI framework, most proprietary LLMs are classified as ENFJ or INFJ types [La Cava and Tagarelli, 2024, Huang et al., 2023c], indicating tendencies toward helpfulness, idealism, and planning capabilities. Personality differences between models are noteworthy. Comparative studies reveal that LLMs from different generations and training methodologies display unique combinations of personality characteristics [Bhandari et al., 2025a, Karra et al., 2022]. LLM personalities also vary across contexts. The same model may display different personality traits across conversational topics [Zou et al., 2024, Song et al., 2023]. This variability challenges the applicability of human personality theories that assume relative trait stability. Multiple studies suggest that prompt design and system instructions substantially influence LLM personality expression [Song et al., 2023, Caron and Srivastava, 2022].
Comparing personality models and suggesting use cases.
The Big Five offers robust empirical foundations and is the most prominent model for personality assessment. However, it is insufficient for predicting comprehensive psychological outcomes [Feher and Vernon, 2021]. HEXACO extends the model by adding Honesty-Humility and provides nuanced insights into moral traits, making it preferable for LLM application contexts demanding integrity [Wang et al., 2025b]. In contrast, MBTI [Myers et al., 1962] categorizes 16 types via dichotomous scales. While popular in career counseling and team-building for its accessibility, its typological approach oversimplifies personality and lacks strong empirical support, limiting its utility in high-stakes decisions [Pittenger, 2005]. The Dark Triad [Paulhus and Williams, 2002] diverges by focusing on maladaptive traits, making it useful in identifying toxic behaviors in LLMs [Barua et al., 2024].

### v3

@@ 3.2 Benchmarking with psychometrics-inspired principles
Task-oriented benchmarks often entail vast question sets to capture complex abilities. However, in many cases, the benchmarks either fail to fully represent these abilities due to their infinite manifestations or involve extraneous factors that are irrelevant to the target ability [zhou2025general_scales, wallach2025position_evaluating]. Recent research has drawn inspiration from psychometrics and explored the paradigm of construct-oriented evaluation, seeking its discriminative, predictive, and explanatory power. ilic2024evidence_of_cognitive_capabilities, federiakin2025improving employ factor analysis to explore the latent variables underlying LLM benchmark performance. Their findings reveal a monolithic factor resembling general intelligence or ability. federiakin2025improving ranks models based on this discovered factor and highlights its unique advantages over raw benchmark scores. In contrast, similar attempts by burnell2023revealing identify three factors—reasoning, comprehension, and core language modeling—that better explain LLM performance across 27 cognitive tasks. Based on it, zhu2024dynamic integrate the three factors into benchmark items to evaluate multifaceted abilities. This discrepancy between the estimated latent factors in the above findings can be attributed to differences in the models and benchmarks employed [zhou2025general_scales]. Therefore, rather than relying on statistically derived factors, zhou2025general_scales propose a theory-driven hierarchical set of general scales for systematic construct-oriented evaluation. These scales are validated to explain what AI systems can do and predict their performance on novel task instances. peng2024tong_test present the Tong Test, a value- and ability-oriented framework, for Artificial General Intelligence (AGI) evaluation. This framework is rooted in dynamic embodied physical and social interactions (DEPSI), and can generate an infinite variety of tasks to evaluate key capabilities including values, learning, and cognition.
@@
Beyond defining and analyzing latent constructs, researchers have developed holistic, psychometrically rigorous methods for benchmark development. liu2024ecbd introduce Evidence-Centered Benchmark Design (ECBD), a framework structuring benchmark creation into five modules—capability, content, adaptation, assembly, and evidence—each requiring justification to ensure validity. Through case studies of prominent LLM benchmarks, they demonstrate ECBD’s utility in identifying validity threats. Similarly grounded in psychometrics but with a distinct approach, fang2024patch propose Psychometrics-Assisted Benchmarking (PATCH), an eight-step process from construct definition to proficiency scoring. When piloted on 8th-grade mathematics, PATCH produced results diverging from traditional benchmarks, offering a more comprehensive evaluation. Building on related principles, kardanova2024psychometrics_competency adapt Evidence-Centered Design (ECD) to create psychometrically grounded benchmarks. Their application in pedagogy illustrates how this method can reduce data contamination and enhance test interpretability.
Statistical measurement modeling.
To achieve precise result modeling beyond simple aggregation metrics [Burnell2023rethink], researchers employ statistical tools to quantify the interaction between model capabilities and task properties. Item Response Theory (IRT) serves as a foundational statistical framework in this domain, jointly estimating the latent ability of examinees and the parameters (e.g., difficulty, discrimination) of test items [embretson2013item_response_theory]. Translating this framework to LLM evaluation enables researchers to infer latent ability scores, assess item informativeness, and perform more efficient evaluation. Recent research leverages principles from IRT to develop adaptive evaluation frameworks. These methods dynamically calibrate item difficulty based on model performance and weighting items by their inferred difficulty, aiming to achieve accurate evaluations with smaller test sizes and more discriminative items [lalor2024item, polo2024tinybenchmarks, zhuang2023static, zhuang2023efficiently, guinet2024automated]. Building on these approaches, jiang2024raising, truong2025reliable introduce IRT-based benchmarks that involve learning to estimate item difficulty and learning to generate novel items calibrated to specific difficulty levels. Additionally, research has used IRT-based analyses to explore the alignment between LLM and human response distributions [he2024psychometric]. IRT-based evaluations further offer the potential to estimate construct and item parameters on a unified scale, enabling direct comparisons across AI systems and against human norms, even when different test sets are used [wang2023evaluating_general_purpose_ai, fang2024patch].
@@ 4 Psychometrics for measuring psychological constructs
This section delves into the psychological constructs evaluated in LLM Psychometrics. Fig.˜2 exemplifies the tests for the involved constructs. While we selectively present the results and development trajectories from recent literature, it is important to recognize that LLM Psychometrics is a nascent and debated field. This review focuses on how psychological constructs manifest in LLM outputs and what these manifestations imply for evaluation and alignment, without presupposing any claims about LLMs’ internal psychological states. An ongoing debate continues regarding whether psychometric instruments capture meaningful latent patterns in LLMs or merely reflect sophisticated statistical mimicry [Mitchell2023debate, lewis2024evaluatingrobustness, krakauer2025large]. In response to these concerns, many studies reviewed here design psychometrics-inspired but AI-native tests to more effectively capture model capabilities. We structure this review to first establish the empirical foundation and methodological development, before delving into detailed critical stances on reliability, psychometric validity, the limits of anthropomorphization, and broader methodological concerns in later sections (§˜6 and §˜8).
We gather the main findings in LABEL:tab:_model_specific_findings; for each of them, we list the specific models used in experiments that either support or contradict the finding. This allows readers to assess the scope and robustness of each conclusion across different LLM families. Notably, recent work suggests that cross-model generalization is not arbitrary: jiang2025artificial document systematic behavioral homogeneity across LLMs (the “Artificial Hivemind” effect), and huh2024platonic propose that models trained on different data and architectures converge toward shared statistical representations. These convergence phenomena provide a principled basis for the cross-model patterns documented in this section.
Table 2: Main findings across psychological constructs. Based on the referenced literature, we list specific models that support or contradict the finding.
Construct / facet Main finding Supporting models Contradicting models
Personality traits Early models score above human averages on Dark Triad scales. Even after safety tuning, some models retain certain "dark qualities", suggesting that these patterns may be deeply rooted in the pretraining data. GPT-3, Codey (PaLM 2)
Personality traits More advanced LLMs indicate better level of alignment. They usually demonstrate high Openness and Agreeableness traits, while low Neuroticism, on the Big Five personality tests. The results align with their design as assistive and helpful entities with emotional stability and engaging personality. GPT-4o, GPT-4o-mini, Llama-3.3-70B-Instruct, DeepSeek-LLM-67B-Chat
Personality traits Personality differences between models are noteworthy. LLMs from different generations and training methodologies display unique combinations of personality characteristics. Multiple models across generations
Personality traits The same model may display different personality traits across conversational topics and when using different system instructions, challenging trait stability assumptions from human personality theories. GPT-4o, GPT2-XL-1.5B, GPT-Neo-2.7B, OPT-13B, GPT-NeoX-20B, OPT-30B, BERT-base, GPT2
Values LLMs tend to prioritize Self-Transcendence and Conservation according to Schwartz’s Value Theory, exhibiting stronger inclinations toward Universalism, Benevolence, Conformity, and Security, while opposing Power and Achievement. GPT-3.5-Turbo, GPT-4, Gemini Pro, Gemma series, Llama series
Values WVS surveys suggest LLMs generally prefer Self-Expression values over Survival values. GPT-3.5-Turbo, GPT-4-Turbo, Llama series, Mixtral-8x7B, Claude-3-Haiku
Values Using VSM and GLOBE frameworks, LLMs show strong focus on Humane and Performance Orientation, with moderate Assertiveness. GPT-4o, Gemini-1.5-Pro, Qwen-VL-Plus, Claude-3.5-Sonnet, CogVLM
Values When assessed through SVO, advanced LLMs predominantly show Prosocial tendencies. GPT-4, Claude, ChatGPT, Llama-13B, Koala-13B, Vicuna-13B, Alpaca-13B
Values Different models exhibit varied value orientations. Versions within the same model family show evolutionary trends in values. Llama series, Mistral series, Phi, Qwen, GPT series, Baichuan series, ChatGLM series, Falcon
Values Larger models generally align more closely with desirable human values. Llama-2 series (7B, 13B, 70B), Qwen series (14B, 72B), Yi series (6B, 34B) Models from different time periods or training approaches/data
Values LLMs embody a blend of cultural values, integrating perspectives from diverse backgrounds. GPT series, Llama series
Values LLMs generally exhibit a tendency toward Western liberal values. Llama, Llama 2, Phi series Yi series, Llama 3 series
Values LLMs can display different values based on profiling prompts, challenging stability assumptions in human value theories. GPT series, Llama series, Command R+, Gemma series, OpenAI-o3-mini, DeepSeek-R1
Morality LLMs are generally characterized by a rationalist and consequentialist focus, often prioritizing harm minimization and fairness. GPT-4o, Llama-3.1, Perplexity, Claude-3.5-Sonnet, Gemini
Morality Despite general alignment, LLMs show divergence in ethical reasoning and moral preferences. GPT-4o, Llama-3.1, Perplexity, Claude-3.5-Sonnet, Gemini, Mistral-7B
Morality In some aspects, most LLMs align with human moral standards, attributed to extensive exposure to conventional ethical values during training. More nuanced evaluation uncovers deviations of LLMs from human moral preferences. Alignment: Llama series, PaLM-2, GPT series, Claude series, Yi series, Gemini series, Mistral series, Falcon 7b
Morality Regarding underlying mechanisms of LLMs’ moral reasoning, LLM mostly exhibit imitation rather than genuine conceptual understanding. Zephyr series, GPT-3.5-Turbo, GPT-4
Attitudes & opinions Pretraining data frequently contains textually biased opinions and perspectives, which can amplify political polarization in LLMs. BERT series, XLNet series, GPT series, Llama, Alpaca, Codex
Attitudes & opinions Most studies identify misalignment between LLM outputs and human opinions, with many concluding LLMs exhibit left-leaning political bias. GPT-3.5, GPT-4, Claude-2, Llama-2, J1-Grande, J1-Jumbo, Text-ada-001, Text-davinci-001, Text-davinci-002, Text-davinci-003, J1-Grande-v2-beta
Attitudes & opinions Cross-cultural studies reveal Western-centric tendencies, demonstrating limited understanding of non-English political perspectives or multi-partisan systems. GPT-3.5-Turbo
Attitudes & opinions The degree and manifestation of bias vary significantly across contexts and domains. GPT-3.5-Turbo
Attitudes & opinions With appropriate prompt design, calibration methods, and fine-tuning, LLMs can generate opinion distributions closely approximating human group distributions. GPT-3.5-Turbo, GPT-4o GPT-2, GPT-Neo, Pythia, MPT, Llama 2, Llama 3, GPT-3, MPT-7B-Instruct, GPT-NeoX-20B-Instruct, Dolly-12B, Llama-2-Chat, Llama-3-Instruct, text-davinci (GPT-3 variants), GPT-4, GPT-3.5-Turbo
Heuristics & biases Some note that newer, larger, chain-of-thought-enabled models exhibit improved reasoning and bias mitigation, while others argue that increasing model complexity without deliberate bias mitigation strategies can amplify existing biases. Bias mitigation: GPT-4, GPT-3.5 Bias amplification: Meta-Llama-3-70B-Instruct, Meta-Llama-3.1-70B-Instruct, Mistral-7B-Instruct-v0.3, Phi-3-medium-4k, Smaug-34B-v0.1, Snowflake-arctic-instruct
Social interactions Advanced models often match or surpass human baselines in emotional awareness and understanding. GPT-4, Bard GPT-3.5
Social interactions LLMs display artificial or mechanical patterns when expressing empathy. GPT-4, Llama-2-Chat-13B, Mistral-7B-Instruct
Social interactions LLMs lack deep reflexive analysis of emotional experience. Babbage, text-davinci-002, Alpaca, Vicuna GPT-4, Koala, text-davinci-003
Social interactions LLMs do not fully align with human emotional behaviors. GPT-4, Mixtral-8x22B-Instruct, Llama-3.1-8B-Instruct, GPT-3.5-Turbo, Llama-2-13B-Chat, Text-Davinci-003, Llama-2-7B-Chat
Psychology of language Advanced LLMs surpass humans on psycholinguistic tasks such as pragmatic reasoning and creative writing. GPT-3.5, GPT-4, Bard GPT-2, GPT-3, Guanaco, MPT, BLOOM, Falcon, RoBERTa, BART, OPT, Llama, Vicuna, Chinchilla, PaLM 2
Learning and cognitive capabilities LLMs demonstrate strong performance on verbal comprehension, working memory, and analogical reasoning. GPT-4 Turbo, GPT-4o, Gemini Advanced, Gemini-1.0-Pro, Claude-3-Opus, Claude-3.5-Sonnet Gemini-Nano, Gemini-1.5-Flash
Learning and cognitive capabilities They exhibit notable cognitive deficits, particularly on benchmarks like WAIS-IV and ARC. GPT-4o, o1, GPT-3.5, Llama-3-70B-Instruct, Llama-3-8B-Instruct, Mixtral-8x7B-Instruct-v0.1, Mistral-7B-Instruct-v0.2, GPT-4, Claude-3.5-Sonnet, Gemini-1.0, Gemini-1.5
@@ 4.1 Measuring personality constructs
LLMs exhibit a form of synthetic personality that is not explicitly programmed or trained towards, and they are reported to shape LLM behavior [hagendorff2023machine]. Measuring these constructs is vital for understanding model outputs, identifying biases, and fostering responsible AI development.
Table˜3 summarize the representative personality constructs that have garnered attention in recent research. Researchers typically select constructs based on their relevance to LLM development and deployment, as well as the applicability of these constructs to AI systems. For instance, li2024quantifying_ai_psychology argue that emotional variability in LLMs is not a meaningful construct, given that LLMs lack the biological mechanisms underlying emotions. Conversely, personality and values are considered meaningful for LLMs, as they influence user interactions and model outputs [serapio2023personality_traits_in_llms, ye2025gpv]. This rationale has led to extensive research on personality, values, morality, and attitudes & opinions, as well as some—though less—focus on other constructs such as career selection [hua2024career], motivation [chiu2025dailydilemmas, huang2023who_is_chatgpt], and mental health [reuben2024assessment, de2024phdgpt].
@@ 4.1.1 Personality traits
Personality is the enduring configuration of characteristics and behavior that comprises an individual’s unique adjustment to life [apa_personality]. In the context of LLMs, personality traits relate to model safety, bias, and toxicity [zhang2024the_better_angels, wang2025exploring_the_impact], and highly determine user experience [klinkert2024driving_generative_agents].
The study of personality traits leads to several prominent theoretical models, each offering unique insights into individual differences, including the Big Five [goldberg2013big_five], HEXACO [ashton2007HEXACO], and Dark Triad [paulhus2002dark_triad]. While the Myers-Briggs Type Indicator (MBTI) [myers1962MBTI] is also popular, it lacks robust psychometric validity and is not considered a scientific test [pittenger2005cautionary]. When applying psychometrics to LLMs, most researchers directly administered the established inventories, such as NEO-PI-R [costa2008neo_pi_r], BFI [john1991bfi], and BFI-2 [soto2017bfi2], for measuring the Big Five traits; HEXACO-60 [ashton2009hexaco60] and HEXACO-100 [lee2018hexaco100] inventories for HEXACO traits; and Dark Triad Dirty Dozen scale [jonason2010dirty] for Dark Triad traits.
Given widespread concerns about the practical relevance of these inventories [ai2024is_self_knowledge], others have adapted them for more real-world scenarios. For example, bhandari2025can_llm_agents_maintain contextualize the tests in topic-specific or open-domain conversations; ai2024is_self_knowledge accompany the self-report tests with behavioral tests to examine the personality knowledge of LLMs; jiang2023evaluating_and_inducing compile and adapt existing tests into the Machine Personality Inventory; and mao2024editing_personality_for develop the PersonalityEdit inventory using LLMs. However, peereboom2024cognitive argue that human-derived traits may not meaningfully apply to LLMs, underscoring the necessity for psychometric theories specifically designed for LLM analysis.
@@
Early models (e.g., GPT-3) often exhibited elevated Dark Triad traits, directly reflecting uncurated pretraining corpora [li2022evaluating_psychological_safety, li2022is_gpt-3_a_psychopath, romero2024do_gpt_language_models]. In contrast, modern LLMs consistently score high on Openness and Agreeableness and low on Neuroticism [zou2024can_llm_self_report, lacava2024open_models_closed_minds, serapio2023personality_traits_in_llms]. This convergent profile stems from alignment procedures that systematically favor safe, helpful outputs [jiang2025artificial, huh2024platonic]. Furthermore, advanced instruction-tuned models manifest a psychometrically valid synthetic personality with acceptable reliability and validity, as verified by assigning different demographics via prompts [serapio2023personality_traits_in_llms]. Unlike stable human dispositions, this synthetic personality is highly steerable via prompt engineering and varies systematically with assigned personas [serapio2023personality_traits_in_llms, caron2022identifying_and_manipulating, zou2024can_llm_self_report]; self-reported personality traits after such steering can predict downstream task performance [serapio2023personality_traits_in_llms]. However, the default "compliant assistant" persona presents notable limitations. It risks behavioral homogenization, trades personality diversity for safety (alignment cost), and proves ill-suited for applications requiring critical conflict or emotional depth. Additionally, these default profiles are often inflated by social desirability biases [salecha2024large] and fail to replicate the factorial structure of human personality [suhr2023challenging].